# Unleashing the Potential of Multi-modal Foundation Models and Video Diffusion for 4D Dynamic Physical Scene Simulation

### v1

@@ Abstract
Realistic simulation of dynamic scenes requires accurately capturing diverse material properties and modeling complex object interactions grounded in physical principles. However, existing methods are constrained to basic material types with limited predictable parameters, making them insufficient to represent the complexity of real-world materials. We introduce a novel approach that leverages multi-modal foundation models and video diffusion to achieve enhanced 4D dynamic scene simulation. Our method utilizes multi-modal models to identify material types and initialize material parameters through image queries, while simultaneously inferring 3D Gaussian splats for detailed scene representation. We further refine these material parameters using video diffusion with a differentiable Material Point Method (MPM) and optical flow guidance rather than render loss or Score Distillation Sampling (SDS) loss. This integrated framework enables accurate prediction and realistic simulation of dynamic interactions in real-world scenarios, advancing both accuracy and flexibility in physics-based simulations.
@@ 1 Introduction
PhysGaussian [40] utilizes 3D Gaussian Splatting along with differentiable MPM to enhance rendering fidelity. Nevertheless, PhysGaussian relies on manual initialization of material properties, limiting its adaptability to complex dynamic scenarios. To enable estimating material properties, recent methods [45, 15, 21] utilize video diffusion to generate videos as motion guidance. However, these methods are limited to elastic material, incapable of accommodating different material types, and restricted to simulating objects exhibiting minor movements. Besides, these methods utilize render loss or Score Distillation Sampling (SDS) [30] to estimate material parameters, which presents challenges: render loss is constrained to small motions due to instability and noise in generated videos, while SDS loss demands high computation memory. In contrast, we investigate that optical flow guidance is both memory-efficient and better suited for capturing large and complex motions, enabling more accurate material parameter optimization (see Tab. 4 and visual comparisons in the Appendix).
PhysGen [22] introduces physics reasoning using large pre-trained visual foundation models, eliminating the need for manual parameter initialization. However, it is restricted to single input images and cannot perform 4D reconstruction or dynamic simulations, limiting its flexibility in simulating complex scenes.
@@
To address these challenges, we propose a novel approach that unleashes the potential of multi-modal foundation models and video diffusion for flexible and high-fidelity 4D dynamic scene simulation (see Fig. 1). A detailed comparison of the feasibility of our approach versus existing methods across various configurations is provided in Tab. 1.
Our method begins by using large pre-trained visual foundation models (e.g., GPT-4   [1]) to identify material types and initialize material parameters based on visual and text-based queries, ensuring that initial simulations are grounded in accurate material properties. Meanwhile, depending on the input type (i.e., multi-view images, dynamic scene videos, or a single image), we employ different scene reconstruction methods [18, 39, 32] to generate 3D Gaussian splats. These splats provide a detailed, differentiable representation of the scene, essential for accurate simulation. The material type and initial material properties inferred from the large visual foundation model are then assigned to the 3D Gaussian splats of the simulating object, ensuring that the simulation starts with realistic physical characteristics. To refine these parameters automatically, we use optical flow as guidance within the video diffusion framework, integrated with a differentiable Material Point Method (MPM). Unlike render loss or SDS loss, which are limited to small and stable motions or computationally intensive, optical flow loss offers a memory-efficient alternative that effectively captures large and complex motions. This enables our method to optimize the material properties to reflect the interactions modeled in the generated video, allowing for simulations that are not only visually plausible but also physically accurate. Our approach supports a diverse range of material behaviors, thus enhancing the flexibility and precision of simulations to represent real-world scenarios comprehensively.
In summary, our contributions are as follows:
@@ 2.1 Dynamic 3D Reconstruction
Dynamic 3D reconstruction has become a fundamental aspect of computer vision, enabling detailed representations of real-world scenes and objects for downstream tasks. Neural Radiance Fields (NeRF) [26] pioneered high-fidelity 3D reconstruction [43, 14, 35, 37, 27, 23] by learning continuous volumetric scene functions from posed images, achieving impressive visual quality. However, NeRF’s high computational demands and reliance on multi-view input limit its effectiveness for dynamic scenes. To enhance real-time reconstruction and efficiency, methods leveraging Gaussian splats [18, 34, 9, 7] for representing surfaces in a differentiable and memory-efficient manner have been developed. Specifically, 3D Gaussian Splatting (3DGS) [18] and PGSR [6] are designed for multi-view images with known poses, enabling high-quality static scene reconstruction. For scenarios involving 4D dynamic scene videos, 4D Gaussian Splatting (4DGS) [39, 5] and Grid4D [17] extend the Gaussian splatting concept to accommodate time-varying data. When only a single image is available, Splatt3R [32] adapts Gaussian splatting to reconstruct scenes without requiring camera parameters or depth information. These Gaussian splatting methods enable continuous representations adaptable to various input types, which is crucial for creating detailed and versatile scene representations, providing the foundation for accurate physics-based simulations in complex environments.
@@ 2.3 Video Generation Models
Video generation models have shown significant potential for simulating dynamic scenes due to their ability to create visually compelling sequences based on learned priors. These models [3, 42, 29, 24, 44] take diverse prompts (e.g., text, image, trajectory, camera control, brush, and etc.) and are trained on extensive datasets to implicitly capture relationships between object appearance and motion.
@@ 3.2 Material Property Initialization
To simulate an object accurately, knowing the material type is essential for applying deformations that adhere to physical laws. Additionally, the initial values of material parameters play a significant role in the subsequent optimization process. Therefore, utilizing foundation models to estimate initial material properties is critical for achieving realistic simulations. In our work, we infer initial material properties by querying GPT-4 [1] with the input image and questions. As shown in Fig. 1, the query involves identifying the closest material type $M$ and inferring its density $\rho$ along with the corresponding initial material parameters ${M_{p}}_{0}$. More details of the material parameters for seven different material types can be found in the Appendix.
@@ 3.3 Material Parameter Optimization with Optical Flow Guidance
$\mathcal{L}_{\text{flow}}=\sum_{t}\|\mathbf{U}(I_{t},I_{t+1})-\hat{\mathbf{U}}(\hat{I}_{t},\hat{I}_{t+1})\|^{2},$ (3)
@@ 4.1 Evaluation on Synthetic Dataset
Metric: We evaluate absolute error (AE) between predicted and ground-truth material parameters to measure the system identification performance across various material types.
Analysis: We report the AE in system identification performance across various material parameters for PAC-NeRF, GIC, and our proposed method on the synthetic dataset in Tab. 2. Our method consistently demonstrates lower AE in most cases, indicating its effectiveness in accurately predicting material properties. Specifically, for Droplet and Letter, our method achieves the lowest absolute errors in viscosity and bulk modulus, demonstrating strong performance in Newtonian fluids. For non-Newtonian fluids like Toothpaste, our approach also shows competitive results, maintaining low shear and bulk modulus errors. For Trophy, representing sand (i.e., granular material), our approach achieves the lowest error in friction angle estimation, highlighting its precision in granular material properties. While system identification remains challenging for complex materials, our method still achieves results comparable to baselines. Besides, as shown in Fig. 2, our method simulates the objects of different material types with deformations closer to the ground truth and sharper textures. More visual comparisons can be found in the Appendix.
Figure 2: Qualitative results of all methods on synthetic dataset.
Object PAC-NeRF [20] GIC [4] Ours
Droplet $\Delta_{\mu}=9$, $\Delta_{\kappa}=0.8\times 10^{4}$ $\Delta_{\mu}=1$, $\Delta_{\kappa}=8.2\times 10^{4}$ $\Delta_{\mu}=\mathbf{0.89}$, $\Delta_{\kappa}=\mathbf{0.3\times 10^{4}}$
Letter $\Delta_{\mu}=16.15$, $\Delta_{\kappa}=3.5\times 10^{4}$ $\Delta_{\mu}=4.95$, $\Delta_{\kappa}=\mathbf{0}$ $\Delta_{\mu}=\mathbf{2.27}$, $\Delta_{\kappa}=0.7\times 10^{4}$
Cream $\Delta_{\mu}=1.110\times 10^{5}$, $\Delta_{\kappa}=5.7\times 10^{5}$, $\Delta_{\mu}=\mathbf{0.003\times 10^{5}}$, $\Delta_{\kappa}=\mathbf{4.8\times 10^{5}}$, $\Delta_{\mu}=0.121\times 10^{5}$, $\Delta_{\kappa}=5.3\times 10^{5}$,
$\Delta_{\tau_{Y}}=0.16\times 10^{3}$, $\Delta_{\eta}=4.4$ $\Delta_{\tau_{Y}}=\mathbf{0.02\times 10^{3}}$, $\Delta_{\eta}=\mathbf{3.4}$ $\Delta_{\tau_{Y}}=1.73\times 10^{3}$, $\Delta_{\eta}=7.5$
Toothpaste $\Delta_{\mu}=1.51\times 10^{3}$, $\Delta_{\kappa}=1.22\times 10^{5}$, $\Delta_{\mu}=0.81\times 10^{3}$, $\Delta_{\kappa}=0.76\times 10^{5}$, $\Delta_{\mu}=\mathbf{0.07\times 10^{3}}$, $\Delta_{\kappa}=\mathbf{0.13\times 10^{5}}$,
$\Delta_{\tau_{Y}}=28$, $\Delta_{\eta}=\mathbf{0.23}$ $\Delta_{\tau_{Y}}=\mathbf{26}$, $\Delta_{\eta}=0.90$ $\Delta_{\tau_{Y}}=46$, $\Delta_{\eta}=28.99$
Torus $\Delta_{E}=4.0\times 10^{4}$, $\Delta_{\nu}=0.022$ $\Delta_{E}=\mathbf{1.0\times 10^{4}}$, $\Delta_{\nu}=\mathbf{0.005}$ $\Delta_{E}=3.9\times 10^{4}$, $\Delta_{\nu}=0.297$
Bird $\Delta_{E}=2.20\times 10^{5}$, $\Delta_{\nu}=0.027$ $\Delta_{E}=\mathbf{0.08\times 10^{5}}$, $\Delta_{\nu}=\mathbf{0.016}$ $\Delta_{E}=0.12\times 10^{5}$, $\Delta_{\nu}=0.171$
Playdoh $\Delta_{E}=1.84\times 10^{6}$, $\Delta_{\nu}=0.028$, $\Delta_{E}=0.42\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.022}$, $\Delta_{E}=\mathbf{0.21\times 10^{6}}$, $\Delta_{\nu}=0.098$,
$\Delta_{\tau_{Y}}=150$ $\Delta_{\tau_{Y}}=\mathbf{20}$ $\Delta_{\tau_{Y}}=41$
Cat $\Delta_{E}=8.39\times 10^{5}$, $\Delta_{\nu}=0.007$, $\Delta_{E}=\mathbf{0.02\times 10^{5}}$, $\Delta_{\nu}=\mathbf{0.004}$, $\Delta_{E}=3.87\times 10^{5}$, $\Delta_{\nu}=0.124$,
$\Delta_{\tau_{Y}}=280$ $\Delta_{\tau_{Y}}=\mathbf{90}$ $\Delta_{\tau_{Y}}=850$
Trophy $\Delta_{\theta_{fric}}=3.9^{\circ}$ $\Delta_{\theta_{fric}}=2.0^{\circ}$ $\Delta_{\theta_{fric}}=\mathbf{0.5^{\circ}}$
Table 2: System identification performance on the synthetic dataset. $\Delta_{*}$ denotes the absolute error (AE) $\downarrow$ for the material parameter *.
@@ 4.2 Evaluation on Real-world Dataset
Baselines: For real-world simulation, we compare our method with the following baselines: PhysGaussian [40], PhysDreamer [45], Physics3D [21]. PhysGaussian employs 3D Gaussian splatting combined with differentiable MPM for high-fidelity rendering and accurate physical simulations but requires manual parameter settings. PhysDreamer and Physcis3D leverage priors from video diffusion models to estimate material properties, but are limited to elastic materials. Note that, PhysGen [22], which also incorporates physical reasoning, is not compared here due to the lack of perceptual information (e.g., albedo) required by PhysGen in the real-world scenarios. Instead, a separate comparison on its synthetic dataset is provided in the Appendix. More details on implementation and minor adaptations can also be found in the Appendix.
Metric: To evaluate performance, we conduct a human evaluation where 32 participants rate 36 videos generated by different methods, presented in random order. Each video is rated on a 5-point Likert scale, from 1 (strongly disagree) to 5 (strongly agree), based on two criteria: physical-realism and photo-realism. Participants independently assess each video along these two dimensions to provide a comprehensive evaluation of simulation realism.
Analysis: To thoroughly evaluate robustness, we applied large input forces in several scenes to induce substantial object deformations and motions, challenging the capacity to accurately simulate intense physical dynamics. As shown in Tab. 3, our method achieves the highest average scores for both physical-realism and photo-realism, significantly outperforming the baseline methods on the real-world dataset. Specifically, our method is the only method achieving scores above 3, while all baseline methods fall below this threshold. The results underscore the limitations of existing methods in capturing complex physical interactions under challenging scenarios.
The qualitative results in Fig. 4 further illustrate the advantages of our method. Even under high-force conditions, our approach produces realistic and stable simulations, capturing both material deformation and dynamic interactions with greater fidelity compared to the baselines. These results highlight the robustness and adaptability of our method for scenes involving diverse materials and large motions, demonstrating the effectiveness of integrating multi-modal foundation models with optical flow-guided optimization. More detailed visual comparisons can be found in the supplementary material.
@@ 4.3 Ablation Study
Effectiveness of Optical Flow Guidance: On the synthetic dataset, we compare the quantitative performance of different loss functions to demonstrate the significance of our optical flow guidance. Specifically, we evaluate the render loss $\mathcal{L}_{render}$, SDS loss (with text prompt) $\mathcal{L}_{sds}$, and our proposed flow loss $\mathcal{L}_{flow}$. As shown in Tab. 4, our optical flow guidance achieves lower Mean Average Errors (MAE) across most material parameters, particularly in viscosity $\eta$, bulk modulus $\kappa$, and yield stress $\tau_{Y}$, indicating more accurate material parameter optimization. The evaluation demonstrates that optical flow guidance significantly improves material parameter optimization and ensures more realistic motion in system identification, compared to the other guidance losses.
Loss MAE $\downarrow$
$\mathcal{L}_{render}$ $\Delta_{\mu}=3.16\times 10^{5}$, $\Delta_{\kappa}=7.03\times 10^{6}$,
$\Delta_{\tau_{Y}}=5.42\times 10^{3}$, $\Delta_{\eta}=2.95\times 10^{2}$,
$\Delta_{E}=1.27\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.15}$, $\Delta_{\theta_{fric}}=29.48^{\circ}$
[…]
$\Delta_{\tau_{Y}}=5.41\times 10^{3}$, $\Delta_{\eta}=2.82\times 10^{2}$,
$\Delta_{E}=1.25\times 10^{6}$, $\Delta_{\nu}=\mathbf{0.15}$, $\Delta_{\theta_{fric}}=29.95^{\circ}$
[…]
$\Delta_{\tau_{Y}}=\mathbf{0.67\times 10^{3}}$, $\Delta_{\eta}=\mathbf{0.18\times 10^{2}}$,
$\Delta_{E}=\mathbf{1.62\times 10^{5}}$, $\Delta_{\nu}=0.17$, $\Delta_{\theta_{fric}}=\mathbf{0.5^{\circ}}$
Table 4: Ablation study of guidance loss on synthetic dataset. We report the Mean Average Errors (MAE) here, and complete results for each object can be found in the Appendix.
Effect of Physics Reasoning with Foundation Model: To evaluate the impact of using a foundation model (e.g., GPT-4) for physics reasoning, we conduct simulations on the real-world dataset with four configurations: (1) manually defined initial material properties without optimization, (2) use manually defined material properties as initialization followed by optimization, (3) use foundation model-inferred properties without optimization, (4) initialize material properties with foundation model-inferred values followed by optimization.
As shown in Fig. 3, simulations initialized with manually defined values produce motions that partially reflect the effects of the input force but lack sufficient fidelity. Further optimization of these manually set values results in overly rigid or tense parameters, leading to less realistic outcomes. In contrast, simulations using foundation model-inferred values without additional optimization provide a more balanced starting point. By initializing with foundation model predictions and optimizing, our method achieves significantly more accurate material dynamics and visually convincing simulations.
@@
Figure 3: Ablation study of different configurations for initializing and optimizing material properties. Each subfigure displays material values, the simulation frame at timestep 30, and space-time slices illustrating deformation frequency.

### v2

@@ Abstract
Realistic simulation of dynamic scenes requires accurately capturing diverse material properties and modeling complex object interactions grounded in physical principles. However, existing methods are constrained to basic material types with limited predictable parameters, making them insufficient to represent the complexity of real-world materials. We introduce PhysFlow, a novel approach that leverages multi-modal foundation models and video diffusion to achieve enhanced 4D dynamic scene simulation. Our method utilizes multi-modal models to identify material types and initialize material parameters through image queries, while simultaneously inferring 3D Gaussian splats for detailed scene representation. We further refine these material parameters using video diffusion with a differentiable Material Point Method (MPM) and optical flow guidance rather than render loss or Score Distillation Sampling (SDS) loss. This integrated framework enables accurate prediction and realistic simulation of dynamic interactions in real-world scenarios, advancing both accuracy and flexibility in physics-based simulations. Our code and data are available at https://zhuomanliu.github.io/PhysFlow
@@ 1 Introduction
PhysGaussian [44] utilizes 3D Gaussian Splatting with differentiable MPM to enhance rendering fidelity, but requires manual initialization of material properties, limiting its adaptability to complex dynamic scenarios. Recent methods [49, 16, 23] use video diffusion for motion guidance to estimate material properties. However, these methods are restricted to elastic material and minor object movements. Besides, these methods utilize render loss [49] or Score Distillation Sampling (SDS) [34] to estimate material parameters, which presents challenges: render loss is constrained to small motions due to instability and noise in generated videos, while SDS loss demands high computation memory. In contrast, we investigate that optical flow guidance is both memory-efficient and better suited for capturing large and complex motions, enabling more accurate material parameter optimization (see Tab. 4 and visual comparisons in the Appendix).
PhysGen [24] introduces physics reasoning using large pre-trained visual foundation models, eliminating the need for manual parameter initialization. However, it is restricted to single input images and simulates objects at a fixed depth and cannot perform 4D reconstruction or dynamic simulations, limiting its flexibility in simulating complex scenes.
@@
To address these challenges, we propose PhysFlow, a novel approach that unleashes the potential of multi-modal foundation models and video diffusion through two key innovations: 1) physics-conditioned flexible material behavior modeling, and 2) optical flow-guided iterative refinement for high-fidelity 4D dynamic scene simulation (see Fig. 1). A detailed comparison of the feasibility of our approach versus existing methods across various configurations is provided in Tab. 1.
Our method leverages large pre-trained visual foundation models (e.g., GPT-4 [1]) to identify material types and initialize material parameters based on visual and text-based queries, ensuring that initial simulations are grounded in accurate material properties. Depending on the input type (i.e., multi-view images, dynamic scene videos, or a single image), we employ different scene reconstruction methods [19, 43, 36] to generate 3D Gaussian splats, providing a differentiable representation for simulation. The material type and initial material properties inferred from the large visual foundation model are then assigned to the 3D Gaussian splats of the simulated object, ensuring that the simulation starts with realistic physical characteristics. To refine these parameters automatically, we use optical flow as guidance within the video diffusion framework, integrated with a differentiable Material Point Method (MPM). Unlike render loss or SDS loss, which are limited to small and stable motions or are computationally intensive, optical flow loss offers a memory-efficient alternative that effectively captures large and complex motions. This enables our method to optimize the material properties to reflect the interactions modeled in the generated video, allowing for simulations that are not only visually plausible but also physically accurate. Our approach supports a diverse range of material behaviors, thus enhancing the flexibility and precision of simulations to represent real-world scenarios comprehensively. In summary, our contributions are as follows:
@@ 2.1 Dynamic 3D Reconstruction
Dynamic 3D reconstruction has become a fundamental aspect of computer vision, enabling detailed representations of real-world scenes and objects for downstream tasks. Neural Radiance Fields (NeRF) [30] pioneered high-fidelity 3D reconstruction [47, 15, 39, 41, 31, 25, 26, 27] by learning continuous volumetric scene functions from posed images, achieving impressive visual quality. However, NeRF’s high computational demands and reliance on multi-view input limit its effectiveness for dynamic scenes. To enhance real-time reconstruction and efficiency, methods leveraging Gaussian splats [19, 38, 9, 7] for representing surfaces in a differentiable and memory-efficient manner have been developed. Specifically, 3D Gaussian Splatting (3DGS) [19] and PGSR [6] are designed for multi-view images with known poses, enabling high-quality static scene reconstruction. For scenarios involving 4D dynamic scene videos, 4D Gaussian Splatting (4DGS) [43, 5] and Grid4D [18] extend the Gaussian splatting concept to accommodate time-varying data. When only a single image is available, Splatt3R [36] adapts Gaussian splatting to reconstruct scenes without requiring camera parameters or depth information. These Gaussian splatting methods enable continuous representations adaptable to various input types, which is crucial for creating detailed and versatile scene representations, providing the foundation for accurate physics-based simulations in complex environments.
@@ 2.3 Video Generation Models
Video generation models have shown significant potential for simulating dynamic scenes due to their ability to create visually compelling sequences based on learned priors. These models [3, 46, 33, 28, 48, 21] take diverse prompts (e.g., text, image, trajectory, camera control, brush, and etc.) and are trained on extensive datasets to implicitly capture relationships between object appearance and motion.
@@ 3.2 Material Property Initialization
To simulate an object accurately, knowing the material type is essential for applying deformations that adhere to physical laws. Additionally, the initial values of material parameters play a significant role in the subsequent optimization process. Therefore, utilizing foundation models to estimate initial material properties is critical for achieving realistic simulations. In our work, we infer initial material properties by querying GPT-4 [1] with the input image and questions. As shown in Fig. 1, the query involves identifying the closest material type $M$ and inferring its density $\rho$ along with the corresponding initial material parameters ${M_{p}}_{0}$. More details of the material parameters for seven different material types can be found in Appendix A.
@@ 3.3 Material Parameter Optimization with Optical Flow Guidance
$\mathcal{L}_{flow}=\sum_{t}\|\mathbf{U}(I_{t},I_{t+1})-\hat{\mathbf{U}}(\hat{I}_{t},\hat{I}_{t+1})\|^{2},$ (3)
@@ 4.1 Evaluation on Synthetic Dataset
Figure 2: Qualitative results of all methods on synthetic dataset.
@@
Metric: We evaluate 1) Relative Error ($\mathbf{RE}=|M_{p}-\hat{M_{p}}|/M_{p}$) comparing predicted versus ground-truth material parameters and 2) End-Point Error (EPE[10] quantifying optical flow deviations to measure the system identification performance across various material types.
Analysis: We report the RE and EPE for system identification performance across various material parameters on the synthetic dataset (Tab. 4) for PAC-NeRF, GIC, and our PhysFlow. Our method consistently achieves lower RE and EPE than PAC-NeRF in most cases and remains competitive with GIC, demonstrating its effectiveness in predicting material properties. Specifically, for Droplet and Letter, our approach yields the lowest RE and EPE, indicating strong performance in Newtonian fluids. For non-Newtonian fluids like Toothpaste, it also delivers competitive results, maintaining low shear and bulk modulus errors. For Trophy, which represents sand (i.e., granular material), our method achieves the lowest error in friction angle estimation, highlighting its precision in granular material properties. While system identification remains challenging for complex materials, our approach performs comparably to baseline methods. As shown in Fig. 2, our method also produces deformations that closely match the ground truth and preserves sharper textures across different material types.
@@ 4.2 Evaluation on Real-world Dataset
PhysGaussian PhysDreamer Physics3D Ours
3.70 3.87 4.41 3.08
[…]
Baselines: For real-world simulation, we compare our method with the following baselines: PhysGaussian [44], PhysDreamer [49], and Physics3D [23]. PhysGaussian combines 3D Gaussian splatting with differentiable MPM for high-fidelity rendering and accurate physical simulations but requires manual parameter settings. PhysDreamer and Physcis3D leverage priors from video diffusion models to estimate material properties, but are limited to elastic materials. To evaluate our method, we apply larger forces than previous baselines. For material settings, we use GPT-queried materials for PhysGaussian, maintain elasticity for PhysDreamer, and incorporate both elasticity and viscoelasticity for Physics3D. PhysGen [24] is not included in this comparison due to its reliance on perceptual information (e.g., albedo), which is unavailable in real-world scenarios. Instead, we provide a separate evaluation on its synthetic dataset, with additional implementation details in Appendix C.3.
Metric: To evaluate performance, we conduct a human evaluation where 32 participants rate 36 videos generated by different methods, presented in random order. Each video is rated on a 5-point Likert scale, from 1 (strongly disagree) to 5 (strongly agree), based on two criteria: physical-realism and photo-realism. Participants independently assess each video along these two dimensions to provide a comprehensive evaluation of simulation realism. Furthermore, to assess motion quality, we consider the Energy Minimization Principle for physically plausible motion and compute the Energy-Constrained Motion Score (ECMS):
$E=\sum_{t}\|\mathbf{v}_{t+1}-\mathbf{v}_{t}\|^{2}+\|\nabla^{2}\mathbf{v}_{t}\|^{2}+\alpha\frac{1}{\sum_{t}\|\mathbf{v}_{t}\|}$ (4)
where $\mathbf{v}_{t}$ denotes velocity and $\alpha=0.1$.
Analysis: To thoroughly evaluate robustness, we applied large input forces in several scenes to induce substantial object deformations and motions, challenging the capacity to accurately simulate intense physical dynamics. As shown in Tab. 2, our method achieves the highest average scores for both physical-realism and photo-realism, significantly outperforming the baseline methods on the real-world dataset. Specifically, our method is the only method achieving scores above 3, while all baseline methods fall below this threshold. And our method achieves the lowest ECMS as shown in Tab. 3, indicating more plausible motion. The results underscore the limitations of existing methods in capturing complex physical interactions under challenging scenarios.
The qualitative results in Fig. 5 further illustrate the advantages of our method. Even under high-force conditions, our approach produces realistic and stable simulations, capturing both material deformation and dynamic interactions with greater fidelity compared to the baselines. These results highlight the robustness and adaptability of our method for scenes involving diverse materials and large motions, demonstrating the effectiveness of integrating multi-modal foundation models with optical flow-guided optimization.
Additional implementation details are provided in Appendix B, and more visual comparisons are available in Appendix C.2.
       Object        PAC-NeRF [22]        GIC [4]        Ours ($\mathcal{L}_{flow}$)        $\mathcal{L}_{render}$        $\mathcal{L}_{sds}$
       Droplet        0.06 / 0.28        0.41 / 0.42        0.02 / 0.08        0.48 / 0.63        0.98 / 1.01
       Letter        0.26 / 0.11        0.02 / 0.07        0.46 / 0.13        0.58 / 0.20        0.70 / 0.18
[…]
       Toothpaste        0.42 / 0.15        0.11 / 0.07        0.24 / 0.16        0.30 / 0.20        0.24 / 0.19
[…]
       Bird        0.08 / 0.62        0.04 / 0.36        0.31 / 0.69        0.37 / 0.88        0.76 / 3.07
[…]
       Cat        0.31 / 0.34        0.02 / 0.08        0.34 / 0.24        0.37 / 0.94        0.85 / 0.70
[…]
       Avg.        0.52 / 0.72        0.11 / 0.38        0.29 / 0.51        0.41 / 1.33        0.70 / 1.90
[…]
w/o GPT Init. GPT Init. w/o Optim. $\mathcal{L}_{render}$ Ours
3.91 3.23 3.21 3.08
[…]
@@ 4.3 Ablation Study
Effectiveness of Optical Flow Guidance: We evaluate the impact of different loss functions on the synthetic dataset to highlight the significance of our optical flow guidance. Specifically, we compare the render loss $\mathcal{L}_{render}=\sum_{t}\lambda L_{1}(I_{t},\hat{I}_{t})+(1-\lambda)L_{D-SSIM}(I_{t},\hat{I}_{t})$ [49], the SDS loss (with text prompt) $\mathcal{L}_{sds}$ [34], and our proposed flow loss $\mathcal{L}_{flow}$. As shown in Tab. 4, optical flow guidance achieves lower RE and EPE across most material parameters, particularly viscosity $\eta$, bulk modulus $\kappa$, and yield stress $\tau_{Y}$ (detailed comparisons for each parameter are provided in Appendix C.4), leading to more accurate material parameter optimization. These results demonstrate that optical flow guidance not only improves parameter estimation but also ensures more realistic motion in system identification compared to other guidance losses. Besides, Fig. 3 and Tab. 5 show that $\mathcal{L}_{flow}$ performs better under large force applications and motions on real-world scenes.
Effect of Physics Reasoning with Foundation Model: To evaluate the impact of using a foundation model (e.g., GPT-4) for physics reasoning, we conduct simulations on the real-world dataset with four configurations: (1) manually defined initial material properties without optimization, (2) use manually defined material properties as initialization followed by optimization, (3) use foundation model-inferred properties without optimization, (4) initialize material properties with foundation model-inferred values followed by optimization. As shown in Fig. 4 and Tab. 5, simulations initialized with manually defined values produce motions that partially reflect the effects of the input force but lack sufficient fidelity. Further optimization of these manually set values results in overly rigid or tense parameters, leading to less realistic outcomes. In contrast, simulations using foundation model-inferred values without additional optimization provide a more balanced starting point. By initializing with foundation model predictions and optimizing, our method achieves significantly more accurate material dynamics and visually convincing simulations.
@@
Figure 4: Ablations on physics reasoning, showing material values, timestep 30 frame, and deformation frequency.